Form Partial differential eq. of u=xnf(y/x)
1
Expert's answer
2017-05-31T12:15:10-0400
du/dx=n*f(y/x)+x*n*(df/dx)*y*(-1)*x^(-2)=n*f(y/x)-(n*(y/x)*(df/dx)).
du/dy=x*n*(df/dy)*x^(-1)=n*(df/dy)